# Density-Aware Farthest Point Sampling

### v1

@@ 4 Bound for the expected prediction error
The optimization problem in (1) cannot be solved directly because the labels of the points in $\mathcal{D}$ and the data source distribution $p_{\mathcal{D}}$ are unknown. To address this challenge, we establish an upper bound for the optimization objective in (1), depending linearly on a weighted fill distance of the training set. Subsequently, we introduce Density-Aware Farthest Point Sampling (DA-FPS), a novel data selection algorithm aiming at selecting training sets that can minimize the proposed bound. First, we define the weighted fill distance, a quantity we can associate with finite subsets of the feature space $\mathcal{X}$ selected from $p_{\mathcal{X}_{\mathcal{L}}}$, which is the marginal on $\mathcal{X}$ of a distribution $p_{\mathcal{L}}\in\mathcal{P}$, called training data distribution.
@@
The weighted fill distance depends on the distance metric and on how we define the weight function $\psi_{\mathcal{L}_{\mathcal{X}}}$. Here, the $L_{2}$-distance is considered, but the following results can be generalized to other distances. Note that if the weight function $\psi_{\mathcal{L}_{\mathcal{X}}}$ is chosen to be constantly equal to one, the weighted fill distance reduces to the standard fill distance. We make a specific choice for the weight function $\psi_{\mathcal{L}_{\mathcal{X}}}$, aiming to take into account the relationship between $p_{\mathcal{X}_{\mathcal{D}}}$ and $p_{\mathcal{X}_{\mathcal{L}}}$. The weight function takes positive or non-negative values in regions of the feature space where $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\leq p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. It takes negative values where $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})>p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. Note that the weighted fill distance is non-negative as we show in appendix˜A. Simply put, the weighted fill distance considers both point distances and distributional differences, and it provides a way to quantify how well the features of points in $\mathcal{L}_{\mathcal{X}}$ selected from $p_{\mathcal{X}_{\mathcal{L}}}$ represent the underlying distribution of data points features arising from $p_{\mathcal{X}_{\mathcal{D}}}$.
@@
For our theoretical analysis, we use assumptions from Climaco & Garcke (2024), which we recapture in Appendix B.1 and B.2. In particular, we consider the assumptions to be valid also for $p_{\mathcal{D}}$ and $p_{\mathcal{L}}$. The first assumption states that given the data feature location ${\bm{X}}={\bm{x}}_{i}$, the associated label value, $y_{i}$, is close to the conditional expected value of the random variable $Y$ at that location. This assumption models scenarios where the true feature-label relationship is stochastic or deterministic but subject to random fluctuations with a magnitude parameterized by a fixed scalar $\epsilon$. Further, a Lipschitz continuity assumption relates to the smoothness of the map connecting $\mathcal{X}$ and $\mathcal{Y}$. It implies that when two data points are closer in $\mathcal{X}$, their corresponding labels tend to be closer in $\mathcal{Y}$. The second assumption essentially states that the expected error on the training set is bounded and that standard regularity assumptions on the error and learned functions hold in the form of Lipschitz continuity. With that, we are able to establish an upper bound for the optimization objective in (1).
@@
$\begin{split}\mathbb{E}_{p_{\mathcal{D}}}&\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]\leq CW_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)+\\ &\quad+\underbrace{\lambda_{l_{\mathcal{Y}}}\epsilon}_{\begin{subarray}{c}\text{labels}\\ \text{uncertainty}\end{subarray}}+\underbrace{\epsilon_{\mathcal{L}}}_{\begin{subarray}{c}\text{max error}\\ \text{training set}\end{subarray}}+\quad\underbrace{\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]}_{\begin{subarray}{c}\text{expected error}\\ \text{training distribution}\end{subarray}}\end{split}$ (4)
@@
where $C:=\left(\lambda_{l_{\mathcal{X}}}+\lambda_{l_{\mathcal{Y}}}\lambda_{p}\right)$. $\lambda_{p}$ Lipschitz constant and $\epsilon$ labels’ uncertainty from Assumption B.1. $\lambda_{l_{\mathcal{X}}}$ and $\lambda_{l_{\mathcal{Y}}}$ are Lipschitz constants of the error function from Assumption B.2. $\epsilon_{\mathcal{L}}$ is the maximum error of $m_{\mathcal{L}}$ on the training set, from Assumption B.2. $W_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)$ weighted fill distance defined as in (2). Moreover, $p_{\mathcal{X}_{\mathcal{D}}}$ and $p_{\mathcal{X}_{\mathcal{L}}}$ are the marginals of $p_{\mathcal{D}}$ and $p_{\mathcal{L}}$, respectively, defined on $\mathcal{X}$.
Theorem 4.2, proof in appendix˜C, provides a qualitative upper bound for the expected value of the error function on the data source distribution $p_{\mathcal{D}}$ depending linearly on the weighted fill distance of the selected training set. Our goal is to design a training data distribution marginal $p_{\mathcal{X}_{\mathcal{L}}}$ and sample a training set $\mathcal{L}_{\mathcal{X}}$ to attempt to minimize the only component of the bound that does not involve the data labels and that we can attempt to estimate before training a regression model: the weighted fill distance.
@@
Note that, the bound in (4) depends on four main quantities: the maximum error on the training set ($\epsilon_{\mathcal{L}}$), which we can only compute after training a regression model, the label uncertainty ($\epsilon$), the expected error on the training data distribution ($\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$), which we may not know or be able to compute, and the weighted fill distance, which contrarily to the previous three quantities does not depend on the data labels. The expected error on the training data distribution and the maximum error on the training set indicate that the bound also depends on how well the unknown trained model fits the training data distribution. We do not consider minimizing these quantities because we can not estimate them without training a regression model. We are considering a passive and model-agnostic sampling scenario where we have no knowledge about the data labels and the regression model at the time of selection. In what follows, we work under the general assumption that the employed trained model works well on the training data and these quantities to be negligible. A similar assumption has been considered in previous work (Just et al., 2023). Note that, under these assumptions, the smaller the weighted fill distance of the selected training set, the smaller the bound for the expected approximation error on data from the source distribution $p_{\mathcal{D}}$. In the next section we propose one possible sampling approach attempting to minimize the weighted fill distance of the selected set.
@@ 5 Density-Aware Farthest Point Sampling (DA-FPS)
In this section we design a data selection procedure aiming at minimizing the weighted fill distance of the selected set: $W_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)$. Unfortunately, given a set $\mathcal{L}_{\mathcal{X}}\subset\mathcal{X}$, its associated weighted fill distance can not be directly computed as the weights explicitly depend on $p_{\mathcal{X}_{\mathcal{D}}}$, and we usually do not know the data source distribution $p_{\mathcal{D}}$, and therefore its marginal, $p_{\mathcal{X}_{\mathcal{D}}}$. Still, we have access to a set $\mathcal{D}_{\mathcal{X}}=\{{\bm{x}}_{i}\}_{i=1}^{n}$ of unlabeled data points arising from $p_{\mathcal{X}_{\mathcal{D}}}$. Thus, we first compute an estimation $\hat{p}_{\mathcal{X}_{\mathcal{D}}}$ of $p_{\mathcal{X}_{\mathcal{D}}}$ from the unlabeled set $\mathcal{D}_{\mathcal{X}}$. Subsequently, we select a training set $\mathcal{L}_{\mathcal{X}}=\{{\bm{x}}_{j}\}_{j=1}^{b}\subset\mathcal{D}_{\mathcal{X}}$ determining an estimation $\hat{p}_{\mathcal{X}_{\mathcal{L}}}$ of the marginal of the training data distribution $p_{\mathcal{X}_{\mathcal{L}}}$. The selection procedure we propose aims at selecting a set $\mathcal{L}_{\mathcal{X}}$ that minimizes a data-driven estimation of the weighted fill distance based on the dataset $\mathcal{D}_{\mathcal{X}}$, the selected subset $\mathcal{L}_{\mathcal{X}}$ and the two estimated marginals $\hat{p}_{\mathcal{X}_{\mathcal{D}}}$ and $\hat{p}_{\mathcal{X}_{\mathcal{L}}}$.

### v2

@@ 4 Bound for the expected prediction error
In this section we develop our main theoretical contribution: an upper bound for the optimization objective in (1) that depends linearly on a novel weighted fill distance of the training set. We first introduce the concept of weighted fill distance of a set. We then establish assumptions and present our main theoretical result.
Since the optimization problem in (1) cannot be solved directly due to the unknown labels of the points in $\mathcal{D}$ and the unknown data source distribution $p_{\mathcal{D}}$, we derive an upper bound for its optimization objective. This bound identifies the weighted fill distance of the training set as a key quantity to minimize for better prediction performance, motivating our Density-Aware Farthest Point Sampling (DA-FPS) algorithm introduced in the next section. We define the weighted fill distance, a quantity we can associate with finite subsets of the feature space $\mathcal{X}$ selected from $p_{\mathcal{X}_{\mathcal{L}}}$, which is the marginal on $\mathcal{X}$ of a distribution $p_{\mathcal{L}}\in\mathcal{P}$, called training data distribution.
@@
The weighted fill distance is a deterministic quantity depending on the densities $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})$ and $p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$, the weight function $\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$, the selected set $\mathcal{L}_{\mathcal{X}}$ and the distance metric. Here, we use the $L_{2}$-metric as it is commonly used in machine learning applications. The following results can be generalized to other metrics. The weighted fill distance is non-negative, as we show in Appendix A, and reduces to the standard fill distance when $\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})=1$ for all ${\bm{x}}\in\mathcal{X}$. Our weight function $\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})=1-\frac{p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})}{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})}$ assigns higher weights to regions where the marginal of training data distribution on $\mathcal{X}$ underrepresents the marginal of the data source distribution, i.e., where $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})<p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. Note that the weighted fill distance can be zero under different circumstances. This occurs, for instance, when $\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})=0$ for all ${\bm{x}}\in\mathcal{X}$, or when any point in the non-zero support of $p_{\mathcal{X}_{\mathcal{D}}}$ is in $\mathcal{L}_{\mathcal{X}}$, that is when $\min\limits_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}=0$ for all $\bm{x}$} such that $p_{\mathcal{D}_{\mathcal{X}}}({\bm{x}})>0$. In the first case, the fact that $\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})=0$ for all ${\bm{x}}\in\mathcal{X}$ implies that the marginals of the training and data source distributions are perfectly aligned, that is, $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})=p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$ for all ${\bm{x}}\in\mathcal{X}$. This simply follows by how we defined the weight function. In the second case, the fact that each point in the non-zero support of $p_{\mathcal{X}_{\mathcal{D}}}$ is in $\mathcal{L}_{\mathcal{X}}$ implies that $\mathcal{L}_{\mathcal{X}}$ already contains all points associated with data locations where a point from the data source distribution could arise. That is, the relevant regions of the data space are already included in the selected set. In other words, the fact that the weighted fill distance is zero can indicate either perfect distributional alignment or full geometric coverage of regions of the data space where points from the data source distribution could arise, but not necessarily both. Thus, the weighted fill distance serves as a joint measure of geometric coverage and distributional matching between the training and data source marginals.
@@
For our theoretical analysis, we use assumptions from Climaco and Garcke (2024), which we recapture in Appendix B.1 and B.2. In particular, we consider the assumptions to be valid also for $p_{\mathcal{D}}$ and $p_{\mathcal{L}}$. The first assumption states that given the data feature location ${\bm{X}}={\bm{x}}_{i}$, the expected associated label value, $y_{i}$, is close to the conditional expected value of the random variable $Y$ at that location. That is, $\mathbb{E}\left[\thinspace|Y-\mathbb{E}[Y|{\bm{x}}_{i}]|\thinspace\big|{\bm{x}}_{i}\right]\leq\epsilon$, with scalar $\epsilon>0$. This assumption models scenarios where the true feature-label relationship is stochastic or deterministic but subject to random fluctuations with a magnitude parameterized by a fixed scalar $\epsilon$. Further, a Lipschitz continuity assumption relates to the smoothness of the map connecting $\mathcal{X}$ and $\mathcal{Y}$. That is, $\bigl|\thinspace\mathbb{E}\left[Y|\hat{{\bm{x}}}\right]-\mathbb{E}\left[Y|\tilde{{\bm{x}}}\right]\thinspace\bigr|\leq\lambda_{p}\|\hat{{\bm{x}}}-\tilde{{\bm{x}}}\|_{2}$, $\forall\;\hat{{\bm{x}}},\tilde{{\bm{x}}}\in\mathcal{X}$, where $\lambda_{p}\in\mathbb{R}^{+}$. It implies that when two data points are closer in $\mathcal{X}$, their corresponding labels tend to be closer in $\mathcal{Y}$. The second assumption essentially states that the expected error on the training set is bounded. That is, we assume there exist $\epsilon_{\mathcal{L}}\geq 0$, depending on the labeled set $\mathcal{L}\subset\mathcal{D}$ and the trained regression model $m_{\mathcal{L}}$, such that for any labeled point $({\bm{x}}_{j},y_{j})\in\mathcal{L}$ we have that $\mathbb{E}[l({\bm{x}}_{j},Y,m_{\mathcal{L}})\big|{\bm{x}}_{j}]\leq\epsilon_{\mathcal{L}}$. We also make standard regularity assumptions on the error and learned functions hold in the form of Lipschitz continuity. With that, we are able to establish an upper bound for the optimization objective in (1).
@@
$\begin{split}\mathbb{E}_{p_{\mathcal{D}}}&\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]\leq CW_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)+\\ &\quad+\Big(\hskip-8.5359pt\underbrace{\lambda_{l_{\mathcal{Y}}}\epsilon}_{\begin{subarray}{c}\text{labels}\\ \text{uncertainty}\end{subarray}}\hskip-2.84544pt+\hskip-2.84544pt\underbrace{\epsilon_{\mathcal{L}}}_{\begin{subarray}{c}\text{max error}\\ \text{training set}\end{subarray}}\hskip-8.5359pt\Big)\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]+\quad\underbrace{\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]}_{\begin{subarray}{c}\text{expected error}\\ \text{training distribution}\end{subarray}},\end{split}$ (4)
@@
where $\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]:=\int_{\mathcal{X}}p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\cdot\mathbf{1}\left\{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})<p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\right\}\,d{\bm{x}}$ quantifies distributional mismatch between $p_{\mathcal{X}_{\mathcal{D}}}$ and $p_{\mathcal{X}_{\mathcal{L}}}$. The indicator function $\mathbf{1}\left\{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})<p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\right\}$ equals 1 when the condition $p_{\mathcal{X}_{\mathcal{D}}}(x)<p_{\mathcal{X}_{\mathcal{L}}}(x)$ is satisfied. $C:=\left(\lambda_{l_{\mathcal{X}}}+\lambda_{l_{\mathcal{Y}}}\lambda_{p}\right)$. $\lambda_{p}$ Lipschitz constant and $\epsilon$ labels’ uncertainty from Assumption B.1. $\lambda_{l_{\mathcal{X}}}$ and $\lambda_{l_{\mathcal{Y}}}$ are Lipschitz constants of the error function from Assumption B.2. $\epsilon_{\mathcal{L}}$ is the maximum error of $m_{\mathcal{L}}$ on the training set, from Assumption B.2. $W_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)$ is the weighted fill distance defined as in (2). Moreover, $p_{\mathcal{X}_{\mathcal{D}}}$ and $p_{\mathcal{X}_{\mathcal{L}}}$ are the marginals of $p_{\mathcal{D}}$ and $p_{\mathcal{L}}$, respectively, defined on $\mathcal{X}$.
Theorem 4.2, proof in Appendix C, provides a qualitative upper bound for the expected value of the error function on the data source distribution $p_{\mathcal{D}}$ depending linearly on the weighted fill distance of the selected training set. Note that if $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})=p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})\;\forall\;{\bm{x}}$ in the non-zero support of $p_{\mathcal{X}_{\mathcal{D}}}$, the weighted fill distance and $\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]$ would be zero and the bound reduces to the trivial but tight inequality $\mathbb{E}_{p_{\mathcal{X}_{\mathcal{D}}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]\leq\mathbb{E}_{p_{\mathcal{D}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$. This shows that when the training and data source distributions are perfectly aligned, our bound reduces to the best achievable bound in this scenario.
@@
The bound in (4) depends on five main quantities: the maximum error on the training set ($\epsilon_{\mathcal{L}}$), which we can only compute after training a regression model, the label uncertainty ($\epsilon$), the expected error on the training data distribution ($\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$), which we may not know or be able to compute, $\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]$ quantifying the mismatch between $p_{\mathcal{X}_{\mathcal{D}}}$ and $p_{\mathcal{X}_{\mathcal{L}}}$, and the weighted fill distance. The last two quantities do not depend on the data labels, contrary to the first three. The expected error on the training data distribution and the maximum error on the training set indicate that the bound also depends on how well the unknown trained model fits the training data distribution. We do not consider minimizing these quantities because we cannot estimate them without training a regression model. We are considering a passive and model-agnostic sampling scenario where we have no knowledge about the data labels and the regression model at the time of selection. In what follows, we work under the general assumption that the employed trained model works well on the training data and that these quantities are negligible. A similar assumption has been considered in previous work (Just et al., 2023). Under such assumptions, the smaller the weighted fill distance of the selected training set and $\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]$, the smaller the bound for the expected approximation error on data from the source distribution $p_{\mathcal{D}}$.
In this work, we focus on minimizing the weighted fill distance of the selected set. We only consider minimizing the weighted fill distance for the following reasons: First, the weighted fill distance provides a quantification of distributional mismatch between the marginals of training and data source distributions by means of the weight function. If there is no distributional mismatch the fill distance is zero. Thus, minimizing the weighted fill distance is related to decreasing $\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]$ as well. Second, by taking into account point distances, the weighted fill distance provides a more comprehensive measure that captures both geometric coverage and distributional alignment aspects. Third, it is more tractable to minimize one of the two quantities than to minimize both. Our goal is to design a data sampling strategy, inducing a training data distribution marginal $p_{\mathcal{X}_{\mathcal{L}}}$, to sample a set $\mathcal{L}_{\mathcal{X}}$ with minimal weighted fill distance.
@@ 5 Density-Aware Farthest Point Sampling (DA-FPS)
In this section we present our main algorithmic contribution: the DA-FPS method for selecting training sets aiming at minimizing the weighted fill distance. We first explain the core algorithmic idea. We then detail how we estimate the weighted fill distance from data. Specifically, we describe the density estimation approach using $k$-nearest neighbors and discuss the estimated weighted fill distance formulation. After that, we present the complete DA-FPS algorithm. The section concludes by addressing computational challenges, providing implementation details, and discussing key algorithmic choices.
DA-FPS is a data selection procedure aiming at minimizing the weighted fill distance of the selected set: $W_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)$. Unfortunately, given a set $\mathcal{L}_{\mathcal{X}}\subset\mathcal{X}$, its associated weighted fill distance cannot be directly computed as the weights explicitly depend on the marginal of the data source distribution $p_{\mathcal{X}_{\mathcal{D}}}$, which we consider to be unknown. Still, we have access to a set $\mathcal{D}_{\mathcal{X}}=\{{\bm{x}}_{i}\}_{i=1}^{n}$ of unlabeled data points arising from $p_{\mathcal{X}_{\mathcal{D}}}$. Thus, we compute an estimation $\hat{p}_{\mathcal{X}_{\mathcal{D}}}$ of $p_{\mathcal{X}_{\mathcal{D}}}$ from the unlabeled set $\mathcal{D}_{\mathcal{X}}$. In this work we assume we do not know a priori the explicit form of the training data distribution marginal $p_{\mathcal{X}_{\mathcal{L}}}$ that we want to use to sample points. If we knew it we could sample according to it and would not have the need to develop a data selection strategy. Instead, we develop a data selection strategy that induces a training data distribution marginal $p_{\mathcal{X}_{\mathcal{L}}}$, which we estimate from the selected $\mathcal{L}_{\mathcal{X}}$ as $\hat{p}_{\mathcal{X}_{\mathcal{L}}}$. The selection procedure we propose aims at selecting a set $\mathcal{L}_{\mathcal{X}}$ that minimizes a data-driven estimation of the weighted fill distance based on the dataset $\mathcal{D}_{\mathcal{X}}$, the selected subset $\mathcal{L}_{\mathcal{X}}$ and the two estimated marginals $\hat{p}_{\mathcal{X}_{\mathcal{D}}}$ and $\hat{p}_{\mathcal{X}_{\mathcal{L}}}$.